a worker earns $6 per hour for the first 40 hours a weeks he works and $9 per hour for any extra hours. assuming that he works at least 40 hours, the total hours worked per week and then simplify?
1
Expert's answer
2013-04-16T11:46:45-0400
A worker works at least 40 hours per week and during that time earns
40*6 = $240
For all extra hours a worker earns:
9*(H - 40) = 9H - 360
Weekly wage worker is the sum of first and second equations:
240 + 9H - 360 = 9H - 120

Answer: So, the worker earns $(9H - 120).